Circular No. 7732

Central Bureau for Astronomical Telegrams INTERNATIONAL ASTRONOMICAL UNION

Mailstop 18, Smithsonian Astrophysical Observatory, Cambridge, MA 02138, U.S.A.

IAUSUBS@CFA.HARVARD.EDU or FAX 617-495-7231 (subscriptions) CBAT@CFA.HARVARD.EDU (science)

        URL http://cfa-www.harvard.edu/iau/cbat.html ISSN 0081-0304

Phone 617-495-7440/7244/7444 (for emergency use only)

SUPERNOVA 2001ev IN UGC 2653 M. Papenkova and W. D. Li, University of California at Berkeley, re-

port the discovery by LOTOSS (cf. IAUC 7514) of an apparent supernova (mag ∼ 18.8) in unfiltered images taken with the Katzman Automatic Imaging Telescope (KAIT) on Oct. 11.5 and 12.3 UT. SN 2001ev is located at α = 3^{h}18^{m}39^{s}.04, δ = +37^{o}36′31′′.8 (equinox 2000.0), which is 8′′.6 east and 2′′.5 north of the nucleus of UGC 2653. KAIT images taken on Sept. 15.5 (limiting mag ∼ 19.8) and 29.5 (limiting mag ∼ 19.0) showed nothing at this position.

SUPERNOVA 2001en IN NGC 523 G. H. Marion, P. Hoeflich, and J. C. Wheeler, University of Texas at

Austin; and W. Vacca, Max-Planck-Institut für Extraterrestrische Physik, write: “SN 2001en (IAUC 7724, 7725) is a type-Ia supernova near maximum light. Near-infrared spectra (range 800–2500 nm) of SN 2001en were obtained on Oct. 8.4 UT using the SpeX instrument at the Infrared Telescope Facility (IRTF). The development of features in the supernova spectra places it temporally between other type-Ia spectra obtained with SpeX at the IRTF: SN 2001br at −2 days (with respect to maximum) and SN 2001dl at +3 days. The spectra of SN 2001en do not show the evolved features shown by type-Ia supernova 2001bg at +10 days. According to W. D. Li, University of California at Berkeley, SN 2001en has brightened from mag 17.0 on Sept. 27.3 to mag 15.0 on Oct. 9.3, and its blue color and brightness are consistent with a type-Ia event in NGC 523.”

T. Matheson, S. Jha, P. Challis, and R. Kirshner, Harvard-Smithsonian Center for Astrophysics, report that a spectrum (range 370–750 nm) of SN 2001en, obtained by J. Huchra on Oct. 11.38 UT with the F. L. Whipple Observatory 1.5-m telescope (+ FAST spectrograph), shows it to be a type-Ia supernova at maximum light. Adopting the NASA/IPAC Extragalactic Database recession velocity of 4758 km/s for the host galaxy, the supernova expansion velocity is ∼ 12 400 km/s for Si II (rest 635.5 nm). The spectralfeature age of the supernova (Riess et al. 1997, A.J. 114, 722) is 0 ± 2 days after maximum light.

A. Dimai and M. Migliardi, Cortina d’Ampezzo, Italy, report that SN 2001en appeared at mag 16.7 on an unfiltered CCD frame (limiting mag 18.5) taken in the course of the CROSS program (cf. IAUC 7373) on Sept. 26.94 UT. Nothing was visible on CROSS images taken on Aug. 27.01 (limiting mag 18.5).

2001 October 12  \begin{matrix}c \\ ◯ \end{matrix} Copyright 2001 CBAT  Daniel W. E. Green